OpenStudy (anonymous):

solve the equation by substitution. 1/10x +1/2y=1/5

OpenStudy (anonymous):

Isolate one variable at a time and use it to fill into the original equation. 1/10x = 1/5 - 1/2y. Multiply both sides by the reciprocal which is 10/1 or 10 to get X by itself. You are left with X = 2 - 5y. Take that equation at fill it into the original. 1/10 (2 - 5y) - 1/2y = 1/5. You get: 1/5 - 1/2y - 1/2y = 1/5. Combine like terms. 1/5 -1y = 1/5. Subtract 1/5 from both sides leaving you -y = 0, multiply by -1 to get rid of the negative sign. Still giving Y = 0. Put 0 in for Y into the original equation. 1/10x + 1/2(0) = 1/5. 1/10x + 0 = 1/5. Subtract 0 giving you 1/10x = 1/5. Multiply both sides by the reciprocal which is 10. X = 2. Your answers are Y = 0, X = 2.